An isoceles triangle has a perimeter of 45cm. Its equal sides each measure 6cm longer than the base. How long is the base of the triangle?: An isoceles triangle has a perimeter of 45cm. Its equal sides each measure 6cm longer than the base. How long is the base of the triangle?
Answer by jojo14344(369)

Let "s"= 2 equal sides, and "b' = base
Since the sides are 6cm longer than the base, "s= b+6"
For the perimeter, we get the measurements of 2 equal sides + the base,
P= 2s + b
45cm = 2(b+6) + b
45 = 2b + 12 + b
45 - 12 = 3b
(33/3) = b ----------- > b= 11cm, base of the triangle
FYI, the sides will be s= 11 + 6 ----> s= 17 cm, 2 equal sides
To check,
P= 2s + b
 45cm = (2*17) + 11
45cm = 34 + 11
45cm = 45cm
Thank you
Jojo